OpenStudy (anonymous):

@ganeshie8 The number of tons of paper waiting to be recycled at a 24 h recycling plant can be modelled by the equation p(t)=0.5 sin((pi/6)t)+4 where t is the time, in hours, and is the number of tons waiting to be recycled. a) Use the equation to estimate the instantaneous rate of change in tons of paper waiting to be recycled when this rate is at its maximum. To make your estimate, use each of the following centred intervals: i) 1 h before to 1 h after the time when the instantaneous rate of change is at its maximum

OpenStudy (anonymous):

How do I determine when the Max Instantaneous rate of change occurs?

OpenStudy (anonymous):

oh i forgot the equation

OpenStudy (anonymous):

its p(t)=0.5 sin((pi/6)t)+4

OpenStudy (anonymous):

Would the Max instantaneous rate of change occur when the slope of the instantaneous rate of change is the steepest?

myininaya (myininaya):

I'm thinking since it says to estimate the instantaneous rate of change we are suppose to find the average rate of change from t-1 to t+1 assuming t is where we have our max instantaneous rate of change then find the max of the resulting function

ganeshie8 (ganeshie8):

yeah i think the teacher wants him to figure out where the max occurs visually and estimate the instantaneous rate of change at that point

ganeshie8 (ganeshie8):

see if this graph helps http://gyazo.com/38e581194095afb7de1665561286990c

OpenStudy (anonymous):

okay so at x=6 is when the instantaneous would approximately equal the max right?

ganeshie8 (ganeshie8):

Yes! thats the first time the instantaneous rate is max after t=0

OpenStudy (anonymous):

okay cool, thanks guys :D

ganeshie8 (ganeshie8):

hey wait, but the graph is decreasing

OpenStudy (anonymous):

okay then i can just choose at x=12 since its increasing?

ganeshie8 (ganeshie8):

Very clever ! yes :)

OpenStudy (anonymous):

Awsome Thanks again for all the help!

ganeshie8 (ganeshie8):

may be just write out that the absolute value of instantaneous rate is maximum at t=6 and t=12 etc.. but the graph is decreasing at t=6, the instantaneous rate of change is negative at t=6 and since the question wants MAXIMUM instantaneous rate of change, it wants positive...

OpenStudy (anonymous):

oh and when i'm calculating rate of change, i choose the point +- 1 hour from the max right?

OpenStudy (anonymous):

so when t=11 and t=13

ganeshie8 (ganeshie8):

yep use t = 12-1 and t = 12+1

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

I got and 0.25 tonnes/ hour, which is right in the book :)

myininaya (myininaya):

just for fun \[\frac{f(t+1)-f(t-1)}{(t+1)-(t-1)} \\ \frac{(.5\sin(\frac{\pi}{6}(t+1))+4)-(.5\sin(\frac{\pi}{6}(t-1))+4)}{2} \\ \frac{.5(\sin(\frac{\pi}{6}t+\frac{\pi}{6})-\sin(\frac{\pi}{6}t-\frac{\pi}{6}))}{2} \\ \frac{1}{4} (\sin(\frac{\pi}{6}t+\frac{\pi}{6})-\sin(\frac{\pi}{6}t-\frac{\pi}{6})) \\ \frac{1}{4}(\sin(\frac{\pi}{6}t)\cos(\frac{\pi}{6})+\sin(\frac{\pi}{6})\cos(\frac{\pi}{6}t)-\sin(\frac{\pi}{6}t)\cos(\frac{\pi}{6})+\sin(\frac{\pi}{6})\cos(\frac{\pi}{6}t)) \\ \frac{1}{4} (\sin(\frac{\pi}{6})\cos(\frac{\pi}{6}t)+\sin(\frac{\pi}{6})\cos(\frac{\pi}{6}t)) \\ \frac{2}{4}\sin(\frac{\pi}{6})\cos(\frac{\pi}{6}t) \\ \frac{1}{2} \sin(\frac{\pi}{6})\cos(\frac{\pi}{6}t) \\ \frac{1}{2} \frac{1}{2} \cos(\frac{\pi}{6}t) \\ \frac{1}{4} \cos(\frac{\pi}{6}t) \] the max of .25cos(pi/6 *t) is .25 to find what what t is when the instanteous rate of change of .25 we do \[\frac{1}{4}\cos(\frac{\pi}{6}t)=\frac{1}{4} \\ \cos(\frac{\pi}{6}t)=1 \\ \frac{\pi}{6}t=2 n \pi \\ t=\frac{6}{\pi}2 n \pi=12n \text{ where } n \text{ is an integer }\] that was just for fun

OpenStudy (anonymous):

O.O no wonder you were taking so long to reply lol

ganeshie8 (ganeshie8):

Wow! that looks a lot better to justify than the graph as the max value of cosine is pretty straightforward to find out xD

myininaya (myininaya):

I thought that is what it wanted us to do when he said some about looking at the interval from an hour before to an hour after

myininaya (myininaya):

but you guys got the answer in like 2 seconds from observation :)

ganeshie8 (ganeshie8):

yeah i feel they want something analytic like that after seeing above method, but it looks like a new method which you invented on the fly hmm never seen/thought about it before but it looks great !

myininaya (myininaya):

well yeah it something I mentioned on the fly and it seemed to work and i thought it could work since it said to estimate the instantaneous rate of change

myininaya (myininaya):

that is a secant line going through the points (t+1,f(t+1)) and (t-1,f(t-1)) where t is assumed to give us the max instantaneous rate change

ganeshie8 (ganeshie8):

I think it works for any interval : \[\frac{f(t+\delta)-f(t-\delta)}{2\delta} \\~\\ \frac{1}{2\delta } \sin (\frac{\pi \delta}{6})\cos(\frac{\pi}{6}t) \]

ganeshie8 (ganeshie8):

max value = \(\large \lim\limits_{\delta \to 0 }\frac{1}{2\delta } \sin (\frac{\pi \delta}{6})\)

ganeshie8 (ganeshie8):

Oh yeah i see where it comes from now, your method is equivalent to average value limit definition of derivative XD

myininaya (myininaya):

https://bell232.wikispaces.com/file/view/MHF4U+N2609.pdf this looks ton easier

myininaya (myininaya):

they just found the period